For the rectangular element considered in Problem 10.52 and Fig. 10.35, find the element nodal force vector due to





an increase in the temperature of the element by 50 C. Perform the needed integration by evaluating the integrand





at the centroid of the element and treating the integrand as a constant throughout the element. Assume a plane stress





condition for the element

1
Expert's answer
2021-12-06T17:19:01-0500

"N_i=\frac{y}{H}- \frac{xy}{LH}"

"N_i=\frac{5}{50}- \frac{6*5}{50+273}"


=0.1 -0.1293

=-0.0293

Ignore negative term

=0.0293